OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7. y = 4x + 10 y = one fourthx + 5 halves y = -4x - 6 y = -one fourthx + 3 halves

OpenStudy (katrinakaif):

parallel = same slope (y-y1) = (m(x-x1) (y- 2) = 4(x +2) Solve

OpenStudy (angela210793):

parallel--->m=m1 y-y1=m(x-x1) y-2=4(x+2) y=4x+10